OpenStudy (anonymous):

Evaluate the integral 7x/ x squared +2 [1,5]

OpenStudy (anonymous):

\[\int\limits_{1}^{5} 7x/ x^2+2 dx\]

OpenStudy (anonymous):

make the substitution \(x^2+2=u\)

OpenStudy (anonymous):

then \[2xdx=du\implies x\,dx={du\over2}\] when \(x=1,u=3\) when \(x=5,u=27\)

OpenStudy (anonymous):

Get the integral of the function and solve for x at 5 and 2

OpenStudy (anonymous):

i mean 5 and 1

OpenStudy (anonymous):

solve the integral in u, replace u with \(x^2+2\) and apply the limits.

OpenStudy (anonymous):

\[ \int_3^{27}{7\over u}{du\over2}={7\over2}[\ln|u|]_3^{27}={7\over2}[\ln(27-\ln(3)]\\ \quad={7\over2}[3\ln(3)-\ln(3)]=\mathbf{7\ln(3)} \]

OpenStudy (anonymous):

I find this way easier